### v2

# Constrained-degree percolation on the hypercubic lattice: uniqueness and some of its consequences

Abstract
We consider the constrained-degree percolation (CDP) model on the hypercubic lattice. This is a continuous-time percolation model defined by a sequence $(U_{e})_{e\in\mathcal{E}^{d}}$ of i.i.d. uniform random variables and a positive integer $k$, referred to as the constraint. The model evolves as follows: each edge $e$ attempts to open at a random time $U_{e}$, independently of all other edges. It succeeds if, at time $U_{e}$, both of its end-vertices have degrees strictly smaller than $k$. It is known [21] that this model undergoes a phase transition when $d\geq 3$ for most nontrivial values of $k$. In this work, we prove that, for any fixed constraint, the number of infinite clusters at any time $t\in[0,1)$ is almost surely either 0 or 1. This uniqueness result implies the continuity of the percolation function in the supercritical regime, $t\in(t_{c},1)$, where $t_{c}$ denotes the percolation critical threshold. The proof relies on a key time-regularity property of the model: the law of the process is continuous with respect to time for local events. In fact, we establish differentiability in time, thereby extending the result of [29] to the CDP setting.
Keywords: constrained percolation; dependent percolation; uniqueness, connectivity function
AMS 1991 subject classification: 60K35; 82C43; 82B43
1 Introduction
The study of random systems with local constraints dates back to the 1930s. One early example is the random sequential adsorption model introduced by Flory [15], motivated by problems in physical chemistry. A few decades later, the classical Bernoulli percolation model was proposed by Broadbent and Hammersley [5] to describe the flow of a deterministic fluid through a random medium. Since then, percolation theory has given rise to a rich variety of models, with applications in fields as diverse as epidemiology [34], composite materials [14], and forest fire dynamics [26].
The intersection of percolation theory and local constraints has attracted attention since at least the work of Gaunt, Guttmann, and Whittington [17], who introduced a percolation model with restricted valence—closely related to the model studied in this paper. In recent years, constrained random structures have gained renewed interest. For instance, Grimmett and Janson [18] studied the Erdős–Rényi random graph $G_{n,p}$ conditioned on vertex degrees lying in a fixed subset $S$ of nonnegative integers. The 1-2 model, in which every vertex on the hexagonal lattice has degree 1 or 2, has been the subject of rigorous mathematical analysis [19]. Kinetically constrained models—particle systems introduced to study glassy and granular materials—have also received considerable attention [3]. Other examples of models with degree or interaction constraints appear in [13, 16, 22, 23].
1.1 The model
Let $\mathbb{L}^{d}=(\mathbb{Z}^{d},\mathcal{E}^{d})$ denote the hypercubic lattice with nearest neighbors. The CDP model we study in this work was introduced in [24, 32]. It is a continuous-time percolation process with infinite-range dependencies, whose dynamics evolve as follows: let $(U_{e})_{e\in\mathcal{E}^{d}}$ be a sequence of $i.i.d.$ random variables with uniform distribution on the interval $[0,1]$ and $k$ be a positive integer, $k\leq 2d$. At time $t=0$, all edges are closed; an edge $e=\langle u,v\rangle$ opens at time $U_{e}$ if, for all $s<U_{e}$, it holds that
$\max\{|\{x\in\mathbb{Z}^{d}:\ \langle x,u\rangle\mbox{ is open at time }s\}|,|\{x\in\mathbb{Z}^{d}:\ \langle x,v\rangle\mbox{ is open at time }s\}|\}<k.$
Once an edge is open, it remains open. We denote by $\mathbb{P}$ the corresponding probability measure on $[0,1]^{\mathcal{E}^{d}}$. Sometimes we refer to $(U_{e})_{e\in\mathcal{E}^{d}}$ as a collection of clocks.
Given a collection of clocks $U=(U_{e})_{e\in\mathcal{E}^{d}}$, let $\omega_{t}^{d,k}(U)\in\{0,1\}^{\mathcal{E}^{d}}$ denote the temporal percolation configuration of open and closed edges at time $t\in[0,1]$. Specifically, $\omega_{t}^{d,k}(U,e)=1$ indicates that $e$ is open at time $t$, while $\omega_{t}^{d,k}(U,e)=0$ indicates that $e$ is closed at time $t$. In this context, we also refer to $e$ as $t$-open or $t$-closed, respectively.
For simplicity, when there is no risk of confusion, we will omit the superscripts $d$ and $k$ and use the notation $\omega_{t}(U)$ and $\omega_{t}(U,e)$ instead.
Consider the CDP model on $\mathbb{L}^{d}$ with $k\leq 2d$. Let $\mathbb{G}_{t}$ be the subgraph of $\mathbb{L}^{d}$ induced by $\omega_{t}$. The connected components of $\mathbb{G}_{t}$ will be called open clusters, and we will say that two sites are connected in $\omega_{t}$, denoted by $x\longleftrightarrow y$, if they are in the same open cluster of $\mathbb{G}_{t}$. If the cluster of a vertice $v$ is infinite, we say that $v$ is connected to infinity, denoted by $v\longleftrightarrow\infty$.
An important object in percolation theory is the probability that the origin belongs to an infinite cluster. With this in mind, define
$\theta^{d,k}(t)=\mathbb{P}\left(0\longleftrightarrow\infty\mbox{ at $t$}\right).$
Clearly, $\theta^{d,k}(t)$ is non-decreasing in $t$. Hence, we define the critical threshold
$t^{d,k}_{c}=\inf\{t\in[0,1]:\ \theta(t)>0\}.$
From now on, we omit the superscripts $d$ and $k$ from our notation and write simply $\theta(t)$ and $t_{c}$ instead of $\theta^{d,k}(t)$ and $t^{k}_{c}(\mathbb{L}^{d})$. Note that for every $k\leq 2d$, the critical threshold $t_{c}$ of the CDP model is always greater than or equal to the critical parameter for Bernoulli percolation on $\mathbb{L}^{d}$. Equality occurs precisely when $k=2d$.
A nice feature of the CDP process is that translations of the lattice are ergodic with respect to $\mathbb{P}$. Since, for all $t\in[0,1]$, the event {there exists an infinite cluster in $\{\mbox{there exists an infinite cluster in }\omega_{t}\}$ is translation invariant, $\theta(t)$ is positive if, and only if,
(there exists an infinite open cluster at $\mathbb{P}\left(\mbox{there exists an infinite open cluster at $t$}\right)=1.$
A non-trivial phase transition for the CDP model on $\mathbb{L}^{2}$ was established in [24], where it was shown that $t_{c}<1$ for $k=3$. Furthermore, the authors of that work prove that if $\theta(t)>0$, then there is a unique infinite open cluster at time $t$, almost surely. Also, when $d>1$ and $k=2$, it was shown that $\theta(t)=0$, for all $t\in[0,1]$.
In [21], the authors provide useful upper bounds for $t_{c}$ for several values of $d$ and $k$. More precisely, they prove that $t_{c}<1.7/d$ when $k>9$ and $d>k/2$; and in the following cases: $(d,k)=(4,7)$; $d\in\{5,6\}$ and $k\geq 8$; $d\in[7,16]$ and $k\geq 9$. In the case $(d,k)=(3,5)$, they show that $t_{c}\leq 0.62$.
1.2 Results
The problem of uniqueness of the infinite open cluster in percolation models goes back to 1960. In that year, Harris [20] showed that Bernoulli percolation on the square lattice admits a unique infinite open cluster in the supercritical regime. Later, Newman and Schulman [25] showed that, under the finite energy condition, the number of infinite open clusters in a percolation process on the $d$-dimensional hypercubic lattice must be 0, 1, or $\infty$, almost surely. The possibility of infinitely many infinite open clusters was excluded in a work by Aizenman, Kesten, and Newman [1]. Burton and Keane [6] gave a beautiful argument for an alternative proof of this last result.
The problem of uniqueness continues to attract the attention of many researchers. In [33], the author shows that, for any translation-invariant Gibbs measure of 1-2 model configurations on the hexagonal lattice, the infinite homogeneous cluster is unique, almost surely. In [9], the authors consider the graph obtained by superposition of an independent pair of uniform infinite non-crossing perfect matchings of the set of integers. They prove that this graph contains at most one infinite path almost surely. Other recent works on the topic include [7, 10, 31].
Remark 1.
The CDP model presents several features that make its analysis particularly challenging. First, it does not satisfy the finite energy condition, which rules out many classical techniques based on local modifications. Second, the model lacks the FKG property, preventing the use of correlation inequalities that are central in standard percolation theory. Finally, it exhibits dependencies of arbitrary order, meaning that local events can be influenced by configurations far away. These aspects make the study of the CDP model both mathematically rich and technically demanding.
The main result of this work is the following.
Theorem 1.
Consider the CDP model on $\mathbb{L}^{d}$ with constraint $k\leq 2d$. Let $N_{t}$ denote the number of infinite clusters at time $t$. Then, for all $t\in[0,1)$, it holds that
$N_{t}\in\{0,1\},\text{ almost surely.}$.
Our second result concerns the connectivity function of two points. Given $x,y\in\mathbb{Z}^{d}$, define
$\tau_{t}(x,y)=\mathbb{P}(x\longleftrightarrow y\mbox{ at time $t$}).$
Theorem 1 allows us to prove the following result.
Theorem 2.
Consider the CDP model on $\mathbb{L}^{d}$ with constraint $k\leq 2d$. If $\theta(t)>0$, then
$\inf_{x,y\in\mathbb{Z}^{d}}\tau_{t}(x,y)>0.$
Clearly, $\theta(t)$ is continuous for $t<t_{c}$, since $\theta(t)=0$ in this case. It turns out that the continuity of $\theta(t)$ for $t>t_{c}$ is a consequence of the uniqueness of the open cluster and of the continuity of $t\mapsto\mathbb{P}(\omega_{t}^{-1}(\cdot))$ for local events. In Section 5, we show that, given any local event $A$, the function $t\mapsto\mathbb{P}(\omega_{t}^{-1}(A))$ is differentiable in the time interval $[0,1]$, improving on Theorem 3 of [29]. This will enable us to prove the following.
Theorem 3.
Consider the CDP model on $\mathbb{L}^{d}$ with constraint $k\leq 2d$. The percolation function $\theta(t)$ is a continuous function of $t$ on the interval $(t_{c},1)$.
Let us now introduce some notation.
1.3 Notation
For $n\in\mathbb{N}$ and $v=(v_{1},v_{2},\ldots,v_{d})\in\mathbb{Z}^{d}$, denote $[n]=\{1,\ldots,n\}$ and $||v||=\sup\{|v_{i}|:i\in[d]\}$. For $A\subset\mathbb{Z}^{2}$, we write the external edge boundary and inner vertex boundary of $A$ as
$\partial^{e}A=\{\langle u,v\rangle\in\mathcal{E}^{d}\colon\ u\in A\mbox{ and }v\in A^{c}\},$
and
$\partial A=\{v\in A:\mbox{ there exits }u\in A^{c}\mbox{ such that }u\sim v\},$
respectively. We denote by
$\mathcal{E}(A)=\{\langle u,v\rangle\in\mathcal{E}^{d}:u,v\in A\}$
the set of edges with both endpoints in $A$. We write $\Lambda_{n}=\{v\in\mathbb{Z}^{d}:||v||\leq n\}$ for the box with side-length $2n$ and centre at the origin.
The restriction of a percolation configuration $\omega$ to a set $K\subset\mathcal{E}^{d}$ is denoted by $\omega|_{K}=\{\omega(e)\colon e\in K\}.$ Define the cylinder set $[\omega]_{K}=\{\omega^{\prime}\colon\omega^{\prime}|_{K}=\omega|_{K}\}.$ We say that the event $A$ lives on a set $K$ if $\omega\in A$ and $\omega^{\prime}\in[\omega]_{K}$ imply $\omega^{\prime}\in A$. Given $K\subset\mathcal{E}^{d}$, we write $\mathcal{H}_{K}=\{A\colon A\mbox{ lives on }K\}$. We say $A$ is a local event if $A$ lives on some finite set $K$.
The remainder of this paper is organized as follows. In Section 2, we introduce the notion of a decreasing cluster. Moreover, we show that, when fixed to a configuration on the exterior edge boundary of a given box, the property of the box being connected to infinity, under certain conditions, depends only on the clocks on the outside of the box (Lemma 1). In Section 3, we present the proof of Theorem 1, in two steps: the first consists of showing that the number of infinite clusters is either 0, 1, or $\infty$, a.s., and the second is to eliminate the case where $N_{t}=\infty$. In Section 4 we prove Theorem 2. Finally, in Section 5 we prove that the function $t\mapsto\mathbb{P}(\omega_{t}^{-1}(\cdot))$ is differentiable on $[0,1]$ for any local event.
2 The decreasing cluster method
As highlighted in Remark 1, a primary challenge of this study is the presence of dependencies of all orders in the CDP model. To address this issue, we focus on analyzing the decreasing cluster of an edge, a concept defined in detail below.
Given a collection of clocks $U\in[0,1]^{\mathcal{E}^{d}}$ and two edges $e$ and $f$, we say that $e$ is bigger than $f$ (with respect to $U$) if $U(e)>U(f)$. A self-avoiding path $P$, consisting of edges $e_{0},e_{1},...,e_{n}$, is called decreasing if, for all $i\in\{1,\dots,n\}$, the edge $e_{i-1}$ is bigger than $e_{i}$.
For $\Lambda\subset\mathcal{E}^{d}$, $e\in\mathcal{E}(\Lambda)$, and $U\in[0,1]^{\mathcal{E}^{d}}$, we define the $\Lambda$-decreasing cluster of $e$ on $U$ as the set $\mathcal{C}^{\Lambda}_{e}(U)$ of edges $f\in\Lambda$ such that there exists a decreasing path, entirely contained in $\Lambda$, starting from $e$ and ending in $f$. For convenience, we assume that $e$ belongs to its decreasing cluster. When $\Lambda=\mathcal{E}^{d}$, we omit $\Lambda$ from the notation and simply write $\mathcal{C}_{e}(U)$. Also, for $\Gamma\subset\mathcal{E}^{d}$, we write
$\mathcal{C}_{\Gamma}(U)=\bigcup_{e\in\Gamma}\mathcal{C}_{e}(U).$ (1)
We stress that the decreasing cluster depends only on the clock variables of the edges and not on their states.
Several authors have used an idea similar to that of the decreasing cluster. For instance, in [2] and [30], this idea is employed to demonstrate that correlations between local events decay at least exponentially fast in the majority percolation and CDP models, respectively. More recently, [8] applied the decreasing cluster framework in the context of the parking process; see also [28]. Additionally, [27] discusses its application in the analysis of random degree models.
Given $\Gamma\subset\mathcal{E}^{d}$, let $\mathcal{F}_{\Gamma}=\{U:|\mathcal{C}_{\Gamma}(U)|<\infty\}$. In what follows, we show that $\mathcal{C}_{\Gamma}$ is almost surely finite.
Proposition 1.
For every $\Gamma\subset\mathcal{E}^{d}$, it holds that $\mathbb{P}(\mathcal{F}_{\Gamma})=1$.
Proof.
For $e\in\mathcal{E}^{d}$ and $n\in\mathbb{N}$, let $A_{n,e}$ be the event that there is a decreasing self-avoiding path of size $n$ starting at $e$. Note that $\{|\mathcal{C}_{e}|=\infty\}\subset A_{n,e}$, for all $n\in\mathbb{N}$. Hence,
$\mathbb{P}(|\mathcal{C}_{e}|=\infty)\leq\frac{2d(2d-1)^{n-1}}{n!},\mbox{ for all }n\in\mathbb{N}.$
Since
$\mathcal{F}_{\Gamma}^{c}=\bigcup_{e\in\Gamma}\{|\mathcal{C}_{e}|=\infty\}.$
the proof is complete. ∎
Remark 2.
From now on, we take $\mathcal{F}$ to be our sample space.
Next, we will present the main result of this section. Before that, we need to introduce further notation. For $E\subset\mathcal{E}^{d}$, let
$\Pi_{E^{c}}:[0,1]^{\mathcal{E}^{d}}\longrightarrow[0,1]^{E^{c}}$
be the canonical projection of $[0,1]^{\mathcal{E}^{d}}$ onto $[0,1]^{E^{c}}$. That is, for each $U\in[0,1]^{\mathcal{E}^{d}}$, $\Pi_{E^{c}}(U)$ is the restriction of $U$ to the set of edges $E^{c}$. We say that $R$ is a modification of $U$ on $E$ if $R\in[0,1]^{E}\times{\Pi_{E^{c}}(U)}$, meaning $R$ agrees with $U$ on all edges outside $E$.
Given $y\in\mathbb{Z}^{d}$, we denote by $\deg_{t}(y,U)$ the number of open edges incident to $y$ in the configuration $\omega_{t}(U)$. In symbols,
$\deg_{t}(y,U)=\#\{e\in\mathcal{E}^{d}\colon\omega_{t}(U,e)=1\mbox{ and }e\sim y\}.$
Given $U\in[0,1]^{\mathcal{E}^{d}}$, let $R$ be a modification of $U$ on $\Lambda_{n}\cup\partial^{e}\Lambda_{n}$. We say that $R$ simplifies the boundary of $\Lambda_{n}$ in $U$ at time $t$ if, for every $e=\langle x,y\rangle\in\partial^{e}\Lambda_{n}$, with $x\in\Lambda_{n}$, it holds that
  1. 1)
    if $\omega_{t}(U,e)=0$, then $R(e)>t$;
  2. 2)
    if $\omega_{t}(U,e)=1$, then $R(e)=U(e)$ and, for all $s<R(e)$, we have $\deg_{s}(x,R)<k$.
It is worth noting the property of state temporal consistency, which ensures that the state of any edge $g$ at time $s:=U(g)$ remains unchanged for all $r>s$. Consequently, if $U(g)=R(g)$, then $\omega_{s}(U,g)=\omega_{s}(R,g)$ if, and only if, $\omega_{r}(U,g)=\omega_{r}(R,g)$ for all $r\geq s$.
The following result establishes that if a modification of a collection of clocks within a box satisfies properties 1) and 2) (as defined above), then the states of the edges outside the box remain identical for both collections of clocks.
Lemma 1.
Let $U\in\mathcal{F}$, and let $R$ be a modification of $U$ in $\mathcal{E}(\Lambda_{n})\cup\partial^{e}\Lambda_{n}$ that simplifies the boundary of $\Lambda_{n}$. Then, for fixed $t\in[0,1)$, it holds that
$\omega_{t}(R,f)=\omega_{t}(U,f),\>\text{ for all }f\in\mathcal{E}(\Lambda_{n})^{c}.$ (2)
Proof.
Let $U\in[0,1]^{\mathcal{E}^{d}}$ as in the statement of the theorem. Fix $t\in[0,1)$ and let $A=\{f\in\partial^{e}\Lambda_{n}:\omega_{t}(U,f)=1\}$. Also, denote $W_{n}=\mathcal{E}(\Lambda_{n})\cup(\partial^{e}\Lambda_{n}-A)$. If $f\in\partial^{e}\Lambda_{n}-A$, then, by the definition of $A$, we have $\omega_{t}(U,f)=0$. Furthermore, since $R$ simplifies the boundary of $\Lambda_{n}$, it follows that $\omega_{t}(R,f)=0$. Therefore, it suffices to verify (2) for $f\in W_{n}^{c}$.
Fix $f\in W_{n}^{c}$. By the definition of $R$, we have $U(f)=R(f)$. Thus, if $U(f)>t$, then $R(f)>t$, and consequently, $\omega_{t}(U,f)=0=\omega_{t}(R,f)$. Therefore, we only need to consider the case where $U(f)\leq t$. Indeed, we will verify a slightly stronger statement. Given $f\in W_{n}^{c}$ with $U(f)\leq t$, we will show (2) for every edge in the decreasing cluster $\mathcal{C}^{\Gamma}_{f}(U)$, where $\Gamma=W_{n}^{c}$, including $f$ itself.
By hypothesis, the decreasing cluster is a.s. finite, so we can write $\mathcal{C}^{\Gamma}_{f}(U)={f_{1},\ldots,f_{m}}$ with $U(f_{1})<\ldots<U(f_{m})=U(f)\leq t$. Consider the following subsets of the decreasing cluster: $F_{0}=\emptyset$ and, for each $i\in[m]$, $F_{i}:=\{f_{1},\dots,f_{i}\}$.
Due to the temporal consistency of states, determining the state of $f_{i}=\langle x_{i},y_{i}\rangle$ requires observing the configuration at time $s=s_{i}:=U(f_{i})\big{(}=R(f_{i})\big{)}$. Denoting $v_{i}\in\{x_{i},y_{i}\}$, we will determine the number of open edges incident to $v_{i}$. Note that
[…]
$=$ $\>\{h\in\mathcal{E}(\Lambda_{n})\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\}$
$\cup\{h\in\partial^{b}\Lambda_{n}-A\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\}$
$\cup\{h\in W_{n}^{c}-\{f_{i}\}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\}.$ (3)
Note that (3) also holds when replacing $U$ with $R$. As seen above, if $h\in\partial^{b}\Lambda_{n}-A$, then $\omega_{t}(U,h)=0=\omega_{t}(R,h)$. Therefore, by the state temporal consistency, we have that, for every $v_{i}\in\{x_{i},y_{i}\}$ and for $T\in\{U,R\}$, it holds that
$\left\{h\in\partial^{b}\Lambda_{n}-A\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(T,h)=1\right\}=\emptyset.$ (4)
Regarding (3), note that, due to the increasing order of elements in $F_{i}$, if $h\in W_{n}^{c}$ such that $h\sim f_{i}$ and $U(h)<U(f_{i})$, we must have $h\in F_{i-1}$. Hence,
$\left\{h\in W_{n}^{c}-\{f_{i}\}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\right\}=\left\{h\in F_{i-1}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\right\}.$ (5)
Therefore, from (3), (4), and (5), we have
$\{h\in\mathcal{E}^{d}-\{f_{i}\}$ $\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\}$
$=$ $\{h\in\mathcal{E}(\Lambda_{n})\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\}$
$\cup\{h\in F_{i-1}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\}.$ (6)
Observe that (5) and (6) still hold when replacing $U$ with $R$. Next, consider $h\in\mathcal{E}(\Lambda_{n})\cup F_{i-1}$. If $h\sim v_{i}$, then we have two possibilities: either $v_{i}\in\partial\Lambda_{n}$ or $v_{i}\notin\Lambda_{n}$.
Case 1. If $v_{i}\in\partial\Lambda_{n}$, we necessarily have $f_{i}\in A$. Therefore, by the definition of $A$, it holds that $\omega_{t}(U,f_{i})=1$. On the other hand, since $R$ simplifies the boundary of $U$, it follows that
$\deg_{t}(v_{i},R)<k.$
That is, the endpoint $v_{i}$ of the edge $f_{i}$ does not prevent it from being open in the configuration $\omega_{t}(R)$.
Case 2. If $v_{i}\notin\Lambda_{n}$, then no edge of $\Lambda_{n}$ is incident to $v_{i}$. Thus, from (6), it follows that
$\left\{h\in\mathcal{E}^{d}-\{f_{i}\}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\right\}=\left\{h\in F_{i-1}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\right\}.$ (7)
Note that (7) still holds when replacing $U$ with $R$. Therefore, to conclude this case, it suffices to verify that for every $i\in[m]$, the following holds:
$\{h\in F_{i-1}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(U,h)=1\}=\{h\in F_{i-1}\colon\ h\sim v_{i}\mbox{ and }\omega_{s}(R,h)=1\},$ (8)
as this, combined with (7), ensures that $\omega_{s}(U,f_{i})=\omega_{s}(R,f_{i})$. This verification can be performed inductively. For $i=1$, we have $F_{i-1}=F_{0}=\emptyset$, so both sets in (8) are empty, resulting in $\omega_{s}(U,f_{1})=\omega_{s}(R,f_{1})$. Now, for $i>1$, if $\omega_{s}(R,f_{j})=\omega_{s}(U,f_{j})$ for $0\leq j<i$, then
$\{g\in F_{i-1}\colon g\sim v_{i},\ \omega_{s}(U,g)=1\}=\{g\in F_{i-1}\colon g\sim v_{i},\ \omega_{s}(R,g)=1\}.$
3 Proof of Theorem 1
In this section we will show Theorem 3. We shall divide the proof in two parts. First, we will show that the number of infinite clusters must almost surely be 0, 1, or $\infty$. This will be addressed in Section 3.1. Next, in Section 3.2, we will rule out the possibility of having infinitely many clusters, thereby completing the proof.
3.1 The number of infinite clusters is either 0, 1, or infinity
Our approach is reminiscent of the work of Newman and Schulman [25]. Specifically, we assume, with positive probability, the existence of $m$ infinite open clusters, where $2\leq m<\infty$. We then select a sufficiently large box that intersects all the existing clusters and modify the configuration within this box in such a way that, with positive probability, all clusters merge into $m-1$ infinite cluster, leading to a contradiction.
This strategy requires careful handling for two key reasons. First, the CDP process does not satisfy the finite-energy property, necessitating a more elaborate construction to replace the original argument in [25]. Second, the model exhibits infinite-range dependencies, meaning that local configuration changes within a box can potentially influence the configuration outside the box.
Let $\mathcal{N}_{t}$ denote the number of infinite open cluster in the temporal configuration $\omega_{t}$, and define the event
$I_{t,m}=\{\mathcal{N}_{t}=m\},\,\,\,\,m=1,\dots,\infty.$
Proposition 2.
For all $t\in[0,1)$, and for every fixed constraint $k\leq 2d$, it holds that
$\mathbb{P}(I_{t,m})=0$
for all $1<m<\infty$.
Proof.
Throughout the proof of Proposition 2, we will rely on a series of claims, the proofs of which are provided in Section 6.
Fix $t\in[0,1)$. Suppose that $\mathbb{P}(I_{t,m})>0$ for some $1<m<\infty$. Since $\mathbb{P}$ is translation-invariant and ergodic, and the event $I_{t,m}$ is also translation-invariant, it follows that $\mathbb{P}(I_{t,m})=1$.
Let $\mathcal{A}_{t,n}=\{\Lambda_{n}\>\text{intersects all the $m$ infinite open clusters at time $t$}\}$. Since $\mathcal{A}_{t,n}\uparrow I_{t,m}$ when $n$ goes to infinity, there exists $n$ large such that the event $\mathcal{A}_{t,n}$ occurs with positive probability. On the event $\mathcal{A}_{t,n}$, we identify $m$ special vertices on $\partial\Lambda_{n}$, each belonging to a distinct infinite cluster. We will employ a strategy to connect two of these vertices via an open path obtained by modifying the configuration inside the box $\Lambda_{n}$, while keeping the external environment intact. This modification produces configurations with $m-1$ infinite clusters. If we show that this happens with positive probability, we obtain a contradiction.
We now focus on the careful selection of points on the boundary of $\Lambda_{n}$ to enable the appropriate modifications within the interior of $\Lambda_{n}$. The selection of boundary vertices is carried out in two stages. Let us discuss the first stage.
We define the set
$F_{n}=\{x\in\partial\Lambda_{n}\colon\#\{y\sim x\colon y\notin\Lambda_{n}\}=1\}$ (9)
as the face of $\Lambda_{n}$. We say that $V=V(m)=\{v_{1},v_{2},\ldots,v_{m}\}\subset F_{n}$ is $m$-connected to infinity if the following conditions are satisfied:
  • a)
    $v_{i}$ is connected to $\infty$ outside $\Lambda_{\bf n-1}$ for every $i\in[m]$;
  • b)
    $v_{i}$ and $v_{j}$ are not connected for $i\neq j$.
Let $\mathcal{X}_{t,n}^{V}$ be the event where $V\subset F_{n}$ is $m$-connected to infinity at time $t$. The next claim shows that $\mathcal{X}_{t,n}^{V}$ occurs with positive probability, for some $V=V(m)\subset F_{n}$.
Claim 1.
Assume that $\mathbb{P}\left(I_{t,m}\right)=1$ for some integer $m>1$. There exist $n\in\mathbb{N}$ and $V=V(m)\subset F_{n}$ such that
$\mathbb{P}\left(\mathcal{X}_{t,n}^{V}\right)>0.$ (10)
Now we discuss the second stage in the choice of the boundary vertices. Let $V=\{v_{1},v_{2},\ldots,v_{m}\}\subset F_{n}$ and $W=\{w_{1},\ldots,w_{m}\}\subset\partial\Lambda_{n}$. We say that $V$ and $W$ are well connected to infinity if
  • c)
    $V$ is $m$-connected to infinity;
  • d)
    $w_{i}$ are $v_{i}$ are connected outside $\Lambda_{\bf n-1}$, for every $i\in[m]$;
  • e)
    $w_{i}$ are connected to infinity outside $\Lambda_{n}$, for every $i\in[m]$.
Let $\mathcal{X}_{t,n}^{V,W}$ be the event that $V$ and $W$ are well connected to infinity at time $t$. It follows from the fact that the number of possible sets with $m$ elements on the boundary of the box is finite, along with (10), that there exists $W\subset\partial\Lambda_{n}$ such that
$\mathbb{P}\left(\mathcal{X}_{t,n}^{V,W}\right)>0.$ (11)
Fix $V$ and $W$ as in (11). Let $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ be a set of $m$ self-avoiding paths in $\Lambda_{n-1}^{c}$ such that $\gamma_{i}$ connects $v_{i}$ to $w_{i}$ for every $i\in[m]$, and consider the event
$\mathcal{Z}_{t}^{\Gamma}:=\{\gamma_{i}\mbox{ is open at time $t$ for all }i\in[m]\}.$
If the number of possible sets $\Gamma$ were finite, we could apply a partition argument similar to the one above to identify a deterministic set $\Gamma$ that connects $V$ and $W$ with positive probability. However, since there are infinitely many possible sets of paths connecting $V$ and $W$, we need a different strategy to limit the number of connections.
Claim 2.
There exists $\Gamma$ such that
$\mathbb{P}\left(\mathcal{X}_{t,n}^{V,W}\cap\mathcal{Z}_{t}^{\Gamma}\right)>0.$ (12)
Our goal now is to adjust the clocks of the edges within the box in such a way that, in the new configuration, at least two infinite components collapse, thereby reducing the total number of such components. To achieve this, we must ensure that the exterior of the box remains unchanged after all modifications. Specifically, we will fix the edge boundary of the box while making appropriate modifications to the interior. For $A\subset\partial^{e}\Lambda_{n}$, define
[…]
In words, on the event $\mathcal{H}_{n,A}^{t}$, an edge $e\in\partial^{e}\Lambda_{n}$ is open at time $t$ if and only if $e\in A$. From (12), it follows that there exists $A\subset\partial^{e}\Lambda_{n}$ such that
$\mathbb{P}\left(\mathcal{X}_{t,n}^{V,W}\cap\mathcal{Z}_{\Gamma}\cap\mathcal{H}_{t,n}^{A}\right)>0.$ (14)
Fix $A$ as in (14) and consider the event
$\mathcal{D}_{t,n}^{\Gamma,A}=\mathcal{X}_{t,n}^{V,W}\cap\mathcal{Z}_{t}^{\Gamma}\cap\mathcal{H}_{t,n}^{A}.$ (15)
Let $U\in\mathcal{D}_{t,n}^{\Gamma,A}$. Suppose there exists a modification $U^{\prime}$ of $U$ on $\mathcal{E}(\Lambda_{n})\cup\partial^{e}\Lambda_{n}$ such that, in $\omega_{t}(U^{\prime})$, the following conditions hold:
  1. A1.
    all edges of $\gamma_{1}$ and $\gamma_{2}$ remain open;
  2. A2.
    all the remaining edges in $\mathcal{E}(\Lambda_{n})$ are closed;
  3. A3.
    the states of all edges in $[\mathcal{E}(\Lambda_{n})]^{c}$ are unchanged and coincide with their states in $\omega_{t}(U)$.
Observe that in $\omega_{t}(U^{\prime})$, the configuration on the exterior of $\Lambda_{n}$ remains identical to $\omega_{t}(U)$, and the paths connecting $v_{1}$ to $w_{1}$ and $v_{2}$ to $w_{2}$ remain open. Therefore, $v_{1}$ and $v_{2}$ remain in infinite components. While closing edges within the box may potentially increase the number of infinite components, this possibility will be disregarded since $N_{t}=m$ is an almost sure event.
Let $\mathcal{U}^{\prime}$ denote the set of clocks $U^{\prime}$ for which there exists $U\in\mathcal{D}_{n,\Gamma,A}^{t}$ such that conditions A1, A2, and A3 occurs. We aim to show that $\mathbb{P}(\mathcal{U}^{\prime})>0$. To this end, define
$G=\mathcal{E}(\Lambda_{n})\cup(\partial^{e}\Lambda_{n}-A),$
and
$\mathcal{E}_{t,n}^{\Gamma,A}=\left\{T\in[0,1]^{G}\colon\bigcap_{\begin{subarray}{c}e\in G\cap(\mathcal{E}(\gamma_{1})\cup\mathcal{E}(\gamma_{2}))\\ f\in G\cap(\mathcal{E}(\gamma_{1})\cup\mathcal{E}(\gamma_{2}))^{c}\end{subarray}}\{T(e)\leq t~{};~{}T(f)>t\}\right\}.$
With a slight abuse of notation, identifying $\mathbb{P}$ with its marginals, and using (14), we obtain
$\mathbb{P}\left(\mathcal{E}_{t,n}^{\Gamma,A}\times\Pi_{G^{c}}\left(\mathcal{D}_{t,n}^{\Gamma,A}\right)\right)=\mathbb{P}\left(\mathcal{E}_{t,n}^{\Gamma,A}\right)\cdot\mathbb{P}\left(\Pi_{G^{c}}\left(\mathcal{D}_{t,n}^{\Gamma,A}\right)\right)>0.$
The next claim shows that $\mathcal{E}_{n,\Gamma,A}^{t}\times\Pi_{G^{c}}\left(\mathcal{D}_{n,\Gamma,A}^{t}\right)\subset\mathcal{U}^{\prime}$.
Claim 3.
Let $U\in\mathcal{D}_{t,n}^{\Gamma,A}$ and let $R\in\mathcal{E}_{t,n}^{\Gamma,A}\times\Pi_{G^{c}}\left(\mathcal{D}_{t,n}^{\Gamma,A}\right)$ be a modification of $U$ on the set $G$. Then,
  • a)
    $R$ simplifies the boundary of $U$. In particular, if $f\in\mathcal{E}(\Lambda_{n})^{c}$, then
    $\omega_{t}(R,f)=\omega_{t}(U,f);$ (16)
  • b)
    the paths $\gamma_{1}$ and $\gamma_{2}$ remain open in $\omega_{t}(R)$.
Denote
$\mathcal{U}^{\prime}=\mathcal{E}_{t,n}^{\Gamma,A}\times\Pi_{G^{c}}\left(\mathcal{D}_{t,n}^{\Gamma,A}\right).$
We now proceed to the final step of the proof, where we connect the infinite clusters of the vertices $v_{1}$ and $v_{2}$. For a given $U\in\mathcal{U}^{\prime}$, let $U^{\prime}$ be a modification of $U$ on $\mathcal{E}(\Lambda_{n-1})\cup\partial^{e}\Lambda_{n-1}$ such that, in $\omega_{t}(U^{\prime})$, the following conditions are satisfied:
  1. B1.
    the states of all edges in $\mathcal{E}(\partial\Lambda_{n}\cup\Lambda_{n}^{c})$ remain identical to their states in $\omega_{t}(U)$;
  2. B2.
    $v_{1}$ and $v_{2}$ are connected in $\omega_{t}(U^{\prime})$.
Let $\mathcal{V}$ denote the set of clocks $U^{\prime}$ for which there exists $U\in\mathcal{U}$ such that conditions B1 and B2 are satisfied. Note that if $T\in\mathcal{V}$, the number of infinite clusters in $\omega_{t}(T)$ will be $m-1$. We will show that $\mathbb{P}(\mathcal{V})>0$.
Since $v_{1}$ and $v_{2}$ are points on the face of $\Lambda_{n}$, there exist $u_{1},u_{2}\in\Lambda_{n-1}$ such that $u_{1}\sim v_{1}$ and $u_{2}\sim v_{2}$. Because $u_{1},u_{2}\in\Lambda_{n-1}$, there exists a self-avoiding path $\alpha$ within $\Lambda_{n-1}$ that connects $u_{1}$ to $u_{2}$. This path $\alpha$ can be empty if $u_{1}=u_{2}$; see Figure 1. Define $\gamma$ as the concatenation
$\gamma=\langle v_{1},u_{1}\rangle\alpha\langle v_{2},u_{2}\rangle.$ (17)
$w_{1}$$w_{2}$$v_{1}$$u_{1}$$v_{2}$$u_{2}$$\Lambda_{n}$$\Lambda_{n-1}$$\gamma_{1}$$\gamma_{2}$$\alpha$
Figure 1: Boxes $\Lambda_{n}$ and $\Lambda_{n-1}$ with the sites $w_{i},v_{i}$ and $u_{i}$, for $i\in\{1,2\}$, highlighted. The paths $\gamma_{1}$ and $\gamma_{2}$ represent the paths that connect $v_{1}$ to $w_{1}$ and $v_{2}$ to $w_{2}$, respectively. The path $\alpha$ is the one that connects $u_{1}$ to $u_{2}$.
Note that $\gamma$ is a self-avoiding path connecting $v_{1}$ to $v_{2}$, with all its edges lying outside $\mathcal{E}(\partial\Lambda_{n}\cup\Lambda_{n}^{c})$. For $H=\partial^{e}\Lambda_{n-1}\cup\mathcal{E}(\Lambda_{n-1})$, define
$\mathcal{E}_{t,n}^{\gamma}=\left\{T\in[0,1]^{H}\colon\bigcap_{\begin{subarray}{c}e\in H\cap\mathcal{E}(\gamma)\\ f\in H-\mathcal{E}(\gamma)\end{subarray}}\{T_{e}\leq t~{};~{}T_{f}>t\}\right\}.$
As before, it follows that
$\mathbb{P}\left(\mathcal{E}_{t,n}^{\gamma}\times\Pi_{H^{c}}\left(\mathcal{U}^{\prime}\right)\right)>0.$
Claim 4.
Let $U\in\mathcal{U}^{\prime}$ and let $R\in\mathcal{E}_{n,\gamma}^{t}\times\Pi_{H^{c}}\left(\mathcal{U}^{\prime}\right)$ be a modification of $U$ on the set $H$. Then,
  • a)
    $v_{1}$ and $v_{2}$ are connected in $\omega_{t}(R)$;
  • b)
    for all $e\in\mathcal{E}(\partial\Lambda_{n}\cup\Lambda_{n}^{c})$, it holds that
    $\omega_{t}(R,e)=\omega_{t}(U,e).$ (18)
Claim 4 establishes that if $R\in\mathcal{E}_{t,n}^{\gamma}\times\Pi_{H^{c}}(\mathcal{U}^{\prime})$, then $\omega_{t}(R)$ satisfy properties B1 and B2 described above for some collection of clocks $U\in\mathcal{U}^{\prime}$. That is, $\mathcal{E}_{t,n}^{\gamma}\times\Pi_{H^{c}}(\mathcal{U}^{\prime})\subset\mathcal{V}$. In particular, $\mathbb{P}(\mathcal{V})>0$.
Observe that if $T\in\mathcal{V}$, the configuration $\omega_{t}(T)$ will contain exactly $m-1$ infinite clusters. Thus, the number of infinite components is $m-1$ with positive probability, leading to a contradiction.
3.2 Impossibility of the existence of an infinite number of infinite clusters
It remains to rule out the possibility that $\mathcal{N}_{t}=\infty$. To do so, we use an argument inspired by the work of Burton and Keane [6], combined with the ideas developed in the previous section.
We say that $v\in\mathbb{Z}^{d}$ is a encounter point for a configuration $\omega\in\{0,1\}^{\mathcal{E}^{d}}$ if the following conditions hold
  • i)
    $v$ belongs to an infinite cluster $C$ in $\omega$;
  • ii)
    the set $C-\{v\}$ has no finite components and exactly three infinite components.
Burton and Keane [6] showed the following geometric result: for any $\omega\in\{0,1\}^{\mathcal{E}^{d}}$ and any rectangle $R\subset\mathbb{Z}^{d}$, the number of encounter points of $\omega$ in $R$ is less than the number of points on the boundary of $R$. Therefore, denoting by $X_{n}$ the random variable that counts the number of encounter points in $\Lambda_{n}$, we have, for all $n\in\mathbb{N}$, that
$X_{n}\leq|\partial\Lambda_{n}|.$ (19)
Denoting by $p$ the probability that the origin is an encounter point, and using the invariance of $\mathbb{P}$ with respect to the translations of the lattice, we obtain, for all $n\in\mathbb{N}$, that
$\mathbb{E}[X_{n}]=p|\Lambda_{n}|,$ (20)
where $\mathbb{E}$ denotes expectation with respect to $\mathbb{P}$.
Assuming the existence of infinitely many infinite clusters, we first establish that $p>0$. Then, using (19), (20), and the amenability of $\mathbb{L}^{d}$, we arrive at a contradiction.
Suppose $\mathbb{P}(N_{t}=\infty)=1$. In this case, the event
$\tilde{\mathcal{A}}_{t,n}:=\{U\in[0,1]^{\mathcal{E}^{d}}:\Lambda_{n}\mbox{ intersects at least 3 infinite clusters in }\omega_{t}(U)\},$
has positive probability for $n$ sufficiently large. For such $n$, we can find $V$, $W$, and $\Gamma$ as in (12), with #$\#V=\#W=3$. Defining $\mathcal{D}_{n,\Gamma,A}^{t}$ as in (15) and following the same steps as in the previous section, we conclude that $\mathbb{P}(\mathcal{D}_{n,\Gamma,A}^{t})>0$. Furthermore, by carefully modifying the clocks of the edges inside $\Lambda_{n}$, while preserving the states of the edges outside $\Lambda_{n}$, we can merge three infinite clusters at a vertex of $\Lambda_{n-1}$. Thus, we prove that some vertex of $\Lambda_{n-1}$ is an encounter point with positive probability. This completes the proof of Theorem 1.
4 Proof of Theorem 2
In this section, we prove Theorem 2, showing that the two-point connectivity function is uniformly bounded away from zero (on the set of pairs of vertices). Here, we write $\mathbb{C}_{t}(x)$ for the open cluster of vertex $x$ at time $t$, i.e., the set of vertices connected to $x$ by a path of $t$-open edges.
Proof of Theorem 2: Given $x\in\mathbb{Z}^{d}$ and $n\in\mathbb{N}$, consider the events
$\mathcal{A}_{t,n}(x)=\{x\longleftrightarrow\partial\Lambda_{n}\mbox{ at $t$},|\mathbb{C}_{t}(x)|<\infty\},$
$\mathcal{B}_{t,n}(x)=\{x\longleftrightarrow\partial\Lambda_{n}\mbox{ at $t$}\},$
and write $\mathcal{B}_{t}(x)=\{x\longleftrightarrow\infty\mbox{ at time $t$}\}.$ Since $\{\mathcal{B}_{t,n}(x)\}_{n\geq 1}$ decreases to $\mathcal{B}_{t}(x)$ we obtain $\mathbb{P}(\mathcal{A}_{t,n}(x))=\mathbb{P}(\mathcal{B}_{t}(x))-\mathbb{P}(\mathcal{B}_{t,n}(x))\rightarrow 0$, when $n$ goes to infinity. Choose $n_{0}$ large enough such that
$\mathbb{P}(\mathcal{A}_{t,n_{0}}(x))\leq\theta^{2}/4.$
This clearly gives $\mathbb{P}(\mathcal{A}_{t,n_{0}}^{c}(x)\cap\mathcal{A}_{t,n_{0}}^{c}(y))\geq 1-\tfrac{1}{2}\theta^{2}$.
Note that $\mathcal{B}_{t}(x)=\mathcal{B}_{t,n}(x)\cap\mathcal{A}_{t,n}^{c}(x)$, for all $n\geq 1$. Hence, by uniqueness of the infinite open cluster, we have
$\tau_{t}(x,y)\geq\mathbb{P}(\mathcal{B}_{t}(x)\cap\mathcal{B}_{t}(y))$ $=\mathbb{P}((\mathcal{B}_{t,n_{0}}(x)\cap\mathcal{A}_{t,n_{0}}^{c}(x))\cap(\mathcal{B}_{t,n_{0}}(y)\cap\mathcal{A}_{t,n_{0}}^{c}(y)))$
$\geq\mathbb{P}(\mathcal{B}_{t,n_{0}}(x)\cap\mathcal{B}_{t,n_{0}}(y))-\tfrac{1}{2}\theta^{2}.$
Since $\mathcal{B}_{t,n_{)}}(x)$ and $\mathcal{B}_{t,n_{0}}(y)$ are local events, Theorem 2 in [29] and translation invariance gives
$\tau_{t}(x,y)\geq\mathbb{P}(\mathcal{B}_{t,n_{0}}(x))\mathbb{P}(\mathcal{B}_{t,n_{0}}(y))-c_{1}n_{0}e^{-\psi d(x,y)}-\tfrac{1}{2}\theta^{2}\geq\tfrac{1}{2}\theta^{2}-c_{1}n_{0}e^{-\psi d(x,y)},$
for some positive constants $\psi$ and $c_{1}$. This gives $\tau_{t}(x,y)>0$ for all $x,y\in\mathbb{Z}^{d}$ with $d(x,y)$ large enough, say $d(x,y)\geq k_{0}$. Since $\tau_{t}(x,y)$ is strictly positive for all $x,y$ with $d(x,y)<k_{0}$, the result follows.
5 Differentiability of $\mathbb{P}$ for local events
In this section, we establish the differentiability of $t\mapsto\mathbb{P}(\omega_{t}^{-1}(\cdot))$ for local events. This result, combined with the uniqueness of the infinite open cluster, will allow us to prove the continuity of the function $\theta(t)$ for $t\in(t_{c}^{k},1)$. Moreover, differentiability of $\mathbb{P}$ plays a key technical role in settings such as dynamic percolation or models where randomness is progressively revealed over time. In particular, when analyzing the sharpness of phase transitions via randomized algorithms on graphs, the differentiability of $\mathbb{P}$ is crucial (see [11, 12]).
Proposition 3.
Let $A$ be a local event. Then, the function $t\mapsto\mathbb{P}_{t}(A)=\mathbb{P}(\omega_{t}^{-1}(A))$ is differentiable on $[0,1]$.
Proof.
Assume that $A$ lives in $\Gamma\subset\mathcal{E}^{d}$. For $t\in[0,1]$ and $e\in\Gamma$, recall the definition in (1) and write
$\mathcal{C}_{t,e}=\left\{\begin{array}[]{cc}\mathcal{C}_{e},&\mbox{ se }U_{e}\leq t,\\ \{e\},&\mbox{ se }U_{e}>t.\\ \end{array}\right.$ (21)
Also, let
$\mathcal{C}_{t}=\bigcup_{e\in\Gamma}\mathcal{C}_{t,e}.$
Denote by $\delta(u,v)$ the graph distance between the vertices $u,v\in\mathbb{Z}^{d}$. Given $\Gamma,\tilde{\Gamma}\subset\mathbb{Z}^{d}$, let $\delta(\Gamma,\tilde{\Gamma})=\inf\{\delta(x,y):x\in\Gamma,y\in\tilde{\Gamma}\}$, and write $B_{r}(\Gamma)=\{x\in\mathbb{Z}^{d}:\delta(x,\Gamma)\leq r\}$ for the ball of radius $r$ centered at $\Gamma$.
Let $r\in\mathbb{N}$, and consider the clock event
$\Xi_{t,r}=\left\{U\colon\mathcal{C}_{t}\subset B_{r}(\Gamma)\right\}.$
A standard counting argument (see [2] and [30]) yields the estimate
$\mathbb{P}\left(\Xi_{t,r}^{c}\right)\leq|\Gamma|\sum_{m=r}^{\infty}\frac{{(2dt)}^{m}}{m!}\leq C(\Gamma,d)\frac{(2dt)^{r}}{r!},$ (22)
with $C=C(\Gamma,d)=|\Gamma|e^{2d}$.
From Proposition 1, we have
$\bigcup_{r=0}^{\infty}\Xi_{t,r}=[0,1]^{\mathbb{E}^{d}}\setminus\mathcal{N},$
with $\mathbb{P}(\mathcal{N})=0$. Write $\Theta_{t,r}=\Xi_{t,r}-\Xi_{t,r-1}$, for $r\geq 1$, and $\Theta_{t,0}=\Xi_{t,0}$. Since $\Xi_{t,r-1}\subset\Xi_{t,r}$ for all $r\geq 1$, we obtain
$\bigcup_{r=0}^{\infty}\Theta_{t,r}=[0,1]^{\mathbb{E}^{d}}\setminus\mathcal{N},$
where the last union is disjoint. Hence,
$\mathbb{P}_{t}(A)\coloneq\mathbb{P}(\omega_{t}^{-1}(A))$ $=\mathbb{P}\left(\bigcup_{r\geq 0}\left(\omega_{t}^{-1}(A)\cap\Theta_{t,r}\right)\right)$
$=\sum_{r\geq 0}\mathbb{P}\left(\omega_{t}^{-1}(A)\cap\Theta_{t,r}\right).$ (23)
A slight modification of Lemma 2 in [29] implies that the clock event $\omega_{t}^{-1}(A)\cap\Xi_{t,r}$ depends only on the random variables $U_{e}$ for $e\in\mathcal{E}(B_{r+1}(\Gamma))$. Hence, for $U,\tilde{U}\in[0,1]^{\mathcal{E}^{d}}$ satisfying $U_{e}=\tilde{U}_{e}$ for all $e$ in this set, $U\in\Xi_{t,r}$ implies $\tilde{U}\in\Xi_{t,r}$. The same holds for $\Theta_{t,r}$ and $\omega_{t}^{-1}(A)\cap\Theta_{t,r}$.
Let $m=m(r)=|\mathcal{E}(B_{r+1}(\Gamma))|$ and write $I_{m}=\{1,2,\ldots,m\}$. Consider the set
[…]
Each element of $\mathcal{P}_{m}$ can be seen as an ordering of the edges of $\mathcal{E}(B_{r+1}(\Gamma))$. For $\sigma\in\mathcal{P}_{m}$ and $\ell\in I_{m}^{*}=I_{m}\cup\{0\}$, define
$\mathbb{E}_{t,r}^{\sigma,\ell}=\{U\in[0,1]^{\mathcal{E}(B_{r+1}(\Gamma))}\colon U_{\sigma(1)}<\ldots<U_{\sigma(\ell)}<t<U_{\sigma(\ell+1)}<\ldots<U_{\sigma(m)}\}.$
A direct calculation shows
$\mathbb{P}\left(\mathbb{E}_{t,r}^{\sigma,\ell}\right)=\frac{t^{\ell}}{\ell!}\cdot\frac{(1-t)^{m-\ell}}{(m-\ell)!}.$
Since $\omega_{t}^{-1}(A)\cap\Theta_{t,r}$ depends only on the clocks in $\mathcal{E}(B_{r+1}(\Gamma))$, we obtain
$\omega_{t}^{-1}(A)\cap\Theta_{t,r}=\bigcup_{\sigma\in\mathcal{P}_{m},\ \ell\in I_{m}^{*}}\left(\omega_{t}^{-1}(A)\cap\Theta_{t,r}\cap\mathbb{E}_{t,r}^{\sigma,\ell}\right).$
A key point is that if $\omega_{t}^{-1}(A)\cap\Theta_{t,r}\cap\mathbb{E}_{t,r}^{\sigma,\ell}\neq\emptyset$, then $\mathbb{E}_{t,r}^{\sigma,\ell}\subset\omega_{t}^{-1}(A)\cap\Theta_{t,r}$. This follows from the fact that the occurrence of the local event $A$ depends only on the order in which clocks ring before and after time $t$, rather than on their actual values. With this in mind, for $r\geq 0$, define
$S_{t,r}=\{(\sigma,\ell)\in\mathcal{P}_{m}\times I_{m}^{*}:\mathbb{E}_{t,r}^{\sigma,\ell}\cap\omega_{t}^{-1}(A)\cap\Theta_{t,r}\neq\emptyset\},$
and note that
$\omega_{t}^{-1}(A)\cap\Theta_{t,r}=\bigcup_{(\sigma,\ell)\in S_{t,r}}\mathbb{E}_{t,r}^{\sigma,\ell}.$
Fix $\ell$ and define $S_{t,r}^{\ell}:=\{\sigma\in\mathcal{P}_{m}:(\sigma,\ell)\in S_{t,r}\}$. Then:
$\mathbb{P}_{t}(A)$ $=\sum_{r=0}^{\infty}\sum_{(\sigma,\ell)\in S_{t,r}}\mathbb{P}(\mathbb{E}_{t,r}^{\sigma,\ell})$
$=\sum_{r=0}^{\infty}\sum_{\ell=0}^{m(r)}\sum_{\sigma\in S_{t,r}^{\ell}}\frac{t^{\ell}}{\ell!}\cdot\frac{(1-t)^{m-\ell}}{(m-\ell)!}.$
Define
$f_{r}(t):=\sum_{\ell=0}^{m(r)}|S_{t,r}^{\ell}|\cdot\frac{t^{\ell}}{\ell!}\cdot\frac{(1-t)^{m-\ell}}{(m-\ell)!}.$ (24)
From (22) and (5), we have
$f_{r}(t)=\mathbb{P}(\omega_{t}^{-1}(A)\cap\Theta_{t,r})\leq\mathbb{P}(\Xi_{t,r-1}^{c})\leq C\cdot\frac{(2dt)^{r-1}}{(r-1)!}.$
Assuming that $|S_{t,r}^{\ell}|$ does not depend on $t$, we obtain (using (24))
$\frac{d}{dt}f_{r}(t)\leq\tfrac{m(r)}{t}\cdot f_{r}(t)\leq Cm(r)\cdot\frac{(2d)^{r-1}}{(r-1)!},$
for all $t\geq 0$, and $r\geq 2$.
Since $m(r)$ is polynomial in $r$, the series $\sum_{r\geq 0}f^{{}^{\prime}}_{r}(t)$ converges uniformly on the interval $[0,1]$ by the Weierstrass M-test, and $\mathbb{P}_{t}(A)$ is differentiable on $[0,1]$.
To complete the proof, it remains to show that $S_{t,r}^{\ell}$ does not depend on $t$. Let $10<t_{1}<t_{2}<$ 1 and fix $r\geq 0$, $\ell\in\{0,\ldots,m\}$. We claim that $S_{t_{1},r}^{\ell}=S_{t_{2},r}^{\ell}$.
Let $\sigma\in S_{t_{1},r}^{\ell}$ and choose $U\in\omega_{t_{1}}^{-1}(A)\cap\Theta_{t_{1},r}\cap\mathbb{E}_{t_{1},r}^{\sigma,\ell}$. Define $V$ by setting $V(e)=U(e)+(t_{2}-t_{1})$ for all $e$. Then:
  • (a)
    $V\in\mathbb{E}_{t_{2},r}^{\sigma,\ell}$,
  • (b)
    $\mathcal{C}_{e}(U)=\mathcal{C}_{e}(V)$ for all $e$, hence $V\in\Theta_{t_{2},r}$,
  • (c)
    since the configuration remains the same, $V\in\omega_{t_{2}}^{-1}(A)$.
Therefore, $V\in\omega_{t_{2}}^{-1}(A)\cap\Theta_{t_{2},r}\cap\mathbb{E}_{t_{2},r}^{\sigma,\ell}$, implying $\sigma\in S_{t_{2},r}^{\ell}$. The reverse inclusion follows similarly by subtracting $(t_{2}-t_{1})$, completing the proof. ∎
Remark 3.
One can show that $t\mapsto\mathbb{P}(\omega_{t}^{-1}(A))$ is infinitely differentiable on $(0,1)$. This follows since the $k$-th derivative $f_{r}^{(k)}(t)$ of $f_{r}(t)$ is bounded by
$f_{r}^{(k)}(t)\leq\left(\frac{m(r)}{t(1-t)}\right)^{k}\cdot f_{r}(t),$
for all $k\in\mathbb{N}$. Hence, for every $\alpha,\beta\in(0,1)$ with $\alpha<\beta$, and $t\in[\alpha,\beta]$, we have $\frac{1}{t}\leq\frac{1}{\alpha}$ and $\frac{1}{1-t}\leq\frac{1}{1-\beta}$, yielding to
$f_{r}^{(k)}(t)\leq\left(\frac{m(r)}{\alpha(1-\beta)}\right)^{k}f_{r}(t).$
The infinite differentiability follows with a second application of the Weierstrass M-test.
We conclude this section with the proof of Theorem 3.
Proof of Theorem 3: Recall the definitions of the events $\mathcal{B}_{t}(0)$ and $\mathcal{B}_{t,n}(0)$ from Section 4. It is clear that
$\theta(t)=\lim_{n\rightarrow\infty}\mathbb{P}(\mathcal{B}_{t,n}(0)).$
By Proposition 3, the function $\mathbb{P}(\mathcal{B}_{t,n})$ is continuous in $t$, so $\theta(t)$ is the decreasing limit of continuous functions. Therefore, $\theta(t)$ is upper semi-continuous. Since $\theta(t)$ is monotonic non-decreasing, it follows that $\theta(t)$ is continuous from the right on $[0,1]$.
Continuity from the left on the interval $(t_{c},1)$ follows from the classical argument of van den Berg and Keane [4], combined with the uniqueness of the infinite open cluster established in Theorem 1. ∎
6 Proofs of claims
In this section, we prove the claims in Section 3.1. We start with the proof of Claim 1.
Proof of Claim 1: Recall that $\mathcal{A}_{t,n}=\{\Lambda_{n}\>\text{intersects all $m$ infinite clusters attime $t$}\}$. If $\mathbb{P}(I_{t,m})=1$, it follows by continuity of the probability that $\mathbb{P}(\mathcal{A}_{n-1})$ goes to 1 as $n$ goes to infinity. On $\mathcal{A}_{n-1}$, we can find $\hat{v}_{1},\ldots,\hat{v}_{m}$ in $\partial\Lambda_{n-1}$ such that
  • a’)
    $\hat{v}_{i}$ is connected to $\infty$ outside $\Lambda_{n-1}$ for every $i\in[m]$;
  • b’)
    $\hat{v}_{i}$ and $\hat{v}_{j}$ are not connected for $i\neq j$.
Condition a’) implies that, for each $i\in[m]$, there exists $v_{i}\in\partial\Lambda_{n}$ such that $\langle\hat{v}_{i},v_{i}\rangle$ is open and $v_{i}$ is connected to $\infty$ outside $\Lambda_{n-1}$. Also, we have that $v_{i}\in F_{n}$ (see (9)). Finally, note that b’) implies b), yielding
$\mathcal{A}_{n-1}\subset\bigcup_{V\subset\Lambda_{n}}\mathcal{X}_{t,n}^{V}.$
Take $n$ sufficiently large and the proof is complete.
[…]
For each $j>n-1$, consider the event
$\mathcal{Y}_{t,j}=\bigcap_{i\in[m]}\left\{v_{i}\overset{\Lambda_{n-1}^{c}\cap\Lambda_{j}}{\leftarrow\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}w_{i}\mbox{ at time $t$}\right\}.$
Note that
$\lim_{j\longrightarrow\infty}\mathbb{P}\left(\mathcal{X}_{t,n}^{V,W}\cap\mathcal{Y}_{t,j}\right)=\mathbb{P}\left(\mathcal{X}_{t,n}^{V,W}\right)>0.$
Fix $j$ large enough such that
$\mathbb{P}(\mathcal{X}_{t,n}^{V,W}\cap\mathcal{Y}_{t,j})>0.$
Let $\mathcal{G}_{j}$ be the collection of sets $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$, where each $\gamma_{i}$ is a self-avoiding path in $\Lambda_{n-1}^{c}\cap\Lambda_{j}$ such that $\gamma_{i}$ connects $v_{i}$ to $w_{i}$ for all $i$ in $[m]$. Note that $\mathcal{G}_{j}$ is finite. Moreover,
$\mathcal{X}_{t,n}^{V,W}\cap\mathcal{Y}_{t,j}=\bigcup_{\Gamma\subset\mathcal{G}_{j}}\left(\mathcal{X}_{t,n}^{V,W}\cap\mathcal{Z}_{t}^{\Gamma}\right),$
from which the desired result follows.
Proof of Claim 3: Let us start with the proof of part (a). Let $e=\langle x,y\rangle\in\partial^{e}\Lambda_{n}$, with $x\in\partial\Lambda_{n}$. By the definition of $\mathcal{E}_{t,n}^{\Gamma,A}$, if $\omega_{t}(U,e)=0$, then $R(e)>t$; whereas if $\omega_{t}(U,e)=1$, then $R(e)=U(e)$. Furthermore, note that any edge $f$ that is closed in $\omega_{t}(U)$ will also be closed in $\omega_{t}(R)$, because in this case $R(f)>t$. Therefore, if $s<R_{e}$, it follows that
$\deg_{s}(x,R)\leq\deg_{s}(x,U)<k.$
Hence, $R$ simplifies the boundary of $\Lambda_{n}$ in $U$ at time $t$. In particular, by Lemma 1, Equation (16) holds.
We now prove (b). By the definition of $\mathcal{E}_{t,n}^{\Gamma,A}$, the only edges in $\mathcal{E}(\Lambda_{n})$ that are open in $\omega_{t}(R)$ are those belonging to the paths $\gamma_{1}$ and $\gamma_{2}$. Conversely, if the paths $\gamma_{1}$ and $\gamma_{2}$ include edges outside of $\mathcal{E}(\Lambda_{n})$, those edges remain open in $\omega_{t}(R)$ due to (16).
Proof of Claim 4 We first prove (a). Let $\gamma$ be as defined in (17). We will show that all edges of $\gamma$ are open in $\omega_{t}(R)$.
First, take an edge $e=\langle u_{i},v_{i}\rangle$ from $\gamma$, $i\in\{1,2\}$. Observe that the number of open edges incident to $u_{i}$ is exactly $2$, specifically the two edges that belong to $\gamma$. On the other hand, the number of open edges incident to $v_{i}$ is at most 3: the edge $e$ itself, the single edge of $\gamma_{i}$ incident to $v_{i}$, and possibly the single edge connecting $v_{i}$ to $\Lambda_{n+1}$ (recall that $v_{i}$ lies on the surface of $\Lambda_{n}$).
Now, suppose $e$ is an edge of the self-avoiding path $\alpha$. In this case, each vertex along $\alpha$ is incident to at most two open edges. Consequently, in $\omega_{t}(R)$, the vertices $v_{1}$ and $v_{2}$ are connected.
For the proof of (b), it is clear that (18) holds for $e\in\mathcal{E}(\partial\Lambda_{n})$. Therefore, it suffices to verify it for $e\in\mathcal{E}(\Lambda_{n})^{c}$. We will check that $R$ simplifies the boundary of $U$.
Suppose $e=\langle x,y\rangle\in\partial^{e}\Lambda_{n}$ with $x\in\Lambda_{n}$. Clearly, $R(e)>t$ whenever $\omega_{t}(U,e)=0$. Conversely, if $\omega_{t}(U,e)=1$, then $R(e)=U(e)$. Also, if $x\not\in\{v_{1},v_{2}\}$, then the number of open edges incident on $x$ in $\omega_{s}(R,e)$ is the same as in $\omega_{s}(U,e)$ for all $s<t$. Hence, for all $s<R(e)$, it holds that
$\deg_{s}(R,e)=\deg_{s}(U,e)<k.$
Finally, if $x\in\{v_{1},v_{2}\}$ and $s<R(e)$, the number of open edges incident to $x$ at time $s$ is at most 2: one connecting the inside and one belonging to $\gamma_{i}$, $i=1,2$.
Acknowledgements
Weberson Arcanjo was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), grant 402952/2023-5. Alan Pereira was partially supported by Fundação de Amparo à Pesquisa do Estado de Alagoas (FAPEAL), process E:60030.0000000161/2022, Project APQ2022021000107, and CNPq, grant 402952/2023-5. Diogo dos Santos was partially supported by CNPq, grants 409198/2021-8 and 402952/2023-5. Roger Silva was partially supported by Fundação de Amparo à Pesquisa do Estado de Minas Gerais (FAPEMIG), grant APQ-06547-24. Marco Ticse was partially supported by FAPEMIG.
Conflict of interest statement
The authors have no conflicts of interest to declare. All co-authors have seen and agree with the contents of the manuscript and there is no financial interest to report. We certify that the submission is original work and is not under review at any other publication.
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.